# Multimodal Emotion Recognition in Conversations: A Survey of Methods, Trends, Challenges and Prospects

### v1

@@ 1 Introduction
However, human emotions are typically conveyed through multiple sensory channels, including hearing, vision, and language. As a result, recent research (e.g., Ma et al., 2024a; Van et al., 2025; Dutta and Ganapathy, 2025) has increasingly focused on multimodal settings in dialogue, a direction we refer to as Multimodal Emotion Recognition in Conversations (MERC). Researchers aim to identify the emotional state of a given utterance by integrating contextual information from different modalities, which often includes subtle personal emotional states such as happiness, anger, and hatred Poria et al. (2019b); Gong et al. (2024), thereby improving the effectiveness of emotion recognition in dialogues. Figure 1 illustrates an example of ERC with textual, acoustic, and visual inputs.
@@ 2 Task Settings and Review Methodology
[…]
Methodology for Literature Compilation.
Strategy. We conduct a comprehensive literature search using sources such as the ACL Anthology, Google Scholar, and general search engines (e.g., Google Chrome). Within the ACL Anthology, we focus on top venues including EMNLP, ACL, NAACL, and related workshops. [^1] We used keywords such as ”multimodal emotion recognition,” ”emotion recognition in conversation,” ”multimodal affective computing,” ”dialogue emotion recognition,” ”MERC benchmark,” ”context-aware emotion detection,” and ”multimodal conversational modeling.”.
@@ 4 Feature Processing
Situation-level. The emotional state of a speaker is influenced not only by the semantic content of the current utterance but also by the surrounding contextual semantics. Therefore, existing approaches Hu et al. (2021a); Majumder et al. (2019) commonly employ specialized networks to model the sequential dependencies among utterances, aiming to more accurately capture the speaker’s temporal emotional dynamics. Given the textual feature $u_{i}\in\mathbb{R}^{d_{u}}$ for each utterance, the sequential context representation $c^{s}_{i}\in\mathbb{R}^{2d_{u}}$ is computed as follows:
@@
Speaker-level. Speaker identity information typically exhibits temporal and relational properties of emotions, which can enhance the model’s ability to perceive speaker role information. Therefore, to more effectively learn and distinguish speaker-level contextual representations, many studies have further introduced speaker-related structured information on top of dialogue context modeling. Common approaches include using Speaker Embeddings Ma et al. (2024a); Shen et al. (2020) to explicitly differentiate between different speakers, or leveraging Graph Neural Networks Ai et al. (2025a); Van et al. (2025) to construct interaction graphs between speakers, thereby more comprehensively modeling the dependencies between them:
  • For Speaker Embeddings: The speaker embedding $S_{i}$ is combined with the modality features (e.g., text, audio, or vision) to produce modality representations that are speaker- and context-aware.
@@
  • For Graph Neural Networks: Speaker interactions can be further modeled by constructing a dialogue graph to capture inter-utterance and inter-speaker dependencies beyond sequential semantics. A dialogue graph is typically defined as $G=(V,E,W,R)$, where each node $v_{i}\in V$ corresponds to the $i$-th utterance, and its associated feature vector $\mathbf{c}_{i}^{s}$ is obtained from sequential modeling of contextual dependencies. Edges $e_{ij}\in E$ represent interaction links between utterances, with associated weights $\omega_{ij}\in W$ reflecting the interaction strength and types $r_{ij}\in R$ encoding speaker-related or structural relationships.
@@ 5.1 Graph-based Methods
Traditional Graph Neural Networks. Previous works such as bc-LSTM Poria et al. (2017) and ICON Hazarika et al. (2018) primarily relied on sequential approaches. DialogueGCN Ghosal et al. (2019) was the first to introduce GNNs into ERC, addressing the limitations of earlier sequence-based models like DialogueRNN Majumder et al. (2019) in capturing contextual dependencies. To effectively integrate information from different modalities, Hu et al. (2021b) constructed a graph structure that fuses multimodal features, enabling the model to capture inter-modal dependencies through graph convolutional networks and incorporating speaker information to enhance the representation of conversational semantics. Inspired by the application of graph convolutions in ERC, Li et al. (2024b) proposed the GSF module, which introduces an alternating graph convolution mechanism to hierarchically extract both cross-modal and intra-modal emotional information. Some studies have further enhanced graph-based models with attention mechanisms for multimodal fusion; for example, Feng and Fan (2025) integrated a Cross-modal Attention Module to better fuse useful information from different modalities, while Nguyen et al. (2023) designed a cross-modal attention mechanism to explicitly model the heterogeneity between modalities.
Hypergraph Neural Networks.
@@
Although traditional graph-based methods can capture long-range and multimodal contextual information, they are often challenged by missing modalities during conversations. Lian et al. (2023) tackled this issue by jointly optimizing classification and reconstruction tasks in an end-to-end manner to effectively model incomplete data. The related works Li et al. (2023b); Huang et al. (2024a) considered the limitations imposed by the pairwise relationships between GNNs nodes.Van et al. (2025) constructed a multimodal fusion graph and introduced Hypergraph Neural Networks Feng et al. (2019) to connect multiple modalities or utterance nodes simultaneously, thereby capturing more complex multivariate dependencies and high-order interactions in conversations, and enhancing the modeling of emotion propagation.
Fourier Graph Neural Networks.
@@
Increasing the depth of GNN layers can lead to the over-smoothing problem Liu et al. (2022); Yi et al. (2023), which hampers the modeling of long-range semantic dependencies and complementary modality relations. To address this issue, GS-MMC Ai et al. (2025b) proposes a graph-based framework for multimodal consistency and complementarity learning. This method employs a Fourier graph operator to extract high- and low-frequency emotional signals from the frequency domain, capturing both local variations and global semantic trends. Additionally, a contrastive learning mechanism Oord et al. (2018) is designed to enhance the semantic consistency and complementarity of these signals in a self-supervised manner, thereby improving the model’s ability to recognize true emotional states.
@@ 5.2 Fusion-based Methods
Equal Modality Weights.
Equal interaction can fully utilize information from various modalities, preventing over-reliance on a single modality. Li et al. (2022) proposed achieving emotion recognition by equally integrating emotional vectors and sentence vectors from different modalities to form emotion capsules. Zhang and Li (2023) designed a local constraint module for modalities within the Transformer to promote modality interaction and incorporates a semantic graph to address the lack of semantic relationship information between utterances. Mao et al. (2021) constructed a hierarchical Transformer where each modality can flexibly switch between sequential and feedforward structures based on contextual information. Inspired by hierarchical modality interaction, Ma et al. (2024a) introduced a hierarchical gating Ma et al. (2019) fusion strategy into the Transformer architecture to enable fine-grained modality interaction and designs self-distillation Zhang et al. (2019) to further learn better modality representations.
Text-Dominant Modality.
Some methods propose models based on primary–auxiliary modality collaboration, where auxiliary modalities are used to enhance the performance of the primary (textual) modality. Huang et al. (2024b) suggested that enhancing a text-dominant model with auxiliary modalities can improve performance. Zou et al. (2022) employed a Transformer architecture to design cross-modal attention for learning fusion relationships between different modalities, preserving the integrity of the primary modality’s features while enhancing the representation of weaker modality features. It also uses a two-stage emotional cue extractor to extract emotional evidence. Building on this, Zou et al. (2023) proposed using weaker modalities as multimodal prompts while performing deep emotional cue extraction for stronger modalities. The cue information is embedded into various attention layers of the Transformer to facilitate the fusion of information between the primary and auxiliary modalities. Zhu et al. (2024) introduced an asymmetric CMA-Transformer module for central and auxiliary modalities to obtain fused modality information and proposes a hierarchical distillation Yang et al. (2021) framework to perform coarse- and fine-grained distillation. This approach ensures the consistency of modality fusion information at different granularities.
@@ 5.3 Generation-based Methods
In recent years, pretrained large language models (LLMs) have achieved remarkable success in natural language processing tasks Chu et al. (2024), demonstrating strong emergent capabilities Wei et al. (2022). However, despite their powerful general-purpose abilities, leveraging their full potential in specific sub-tasks still requires carefully crafted, high-quality prompts Wei et al. (2021) to bridge the gap in reasoning capabilities.As shown in Figure 5, researchers have proposed various model improvement strategies to effectively integrate contextual and multimodal information into LLMs while addressing their substantial computational resource demands.
@@
Instruction-Tuned with Speaker and Context Modeling.
ERC tasks have predominantly relied on discriminative modeling frameworks. With the emergence of LLMs, InstructERC Lei et al. (2023) was the first to propose a generative framework for ERC. It introduces a simple yet effective retrieval-based prompting module that helps LLMs explicitly integrate multi-granularity supervisory signals from dialogues. Additionally, it incorporates an auxiliary emotion alignment task to better model the complex emotional transitions between interlocutors in conversations. Inspired by the integration of commonsense knowledge in COSMIC Ghosal et al. (2020), recent work Fu (2024); Fu et al. (2025a) designed a prompt generation approach based on dialogue history to elicit speaker-related commonsense using LLMs by injecting commonsense knowledge into ERC.
Behavior-Aware and Multimodal Instruction-Tuned.
To address the lack of multimodal integration, Dutta and Ganapathy (2025) incorporated both acoustic and textual modalities. Considering that visual information may provide richer emotional cues, Zhang et al. (2024b) constructed a high-quality instruction dataset using image and text data, and fine-tunes the model using Low-Rank Adaptation (LoRA) Song et al. (2024). Furthermore, Fu et al. (2025b) introduced a novel behavior-aware Multimodal LLM (MLLM)-based ERC framework. It consists of three core components: a video-derived behavior generation module, a behavior alignment and refinement module, and an instruction tuning module Wei et al. (2021). The first two modules enable the model to infer human behaviors from limited information, thereby enhancing its behavioral perception capability. The instruction tuning module improves the model’s emotion recognition performance by aligning and fine-tuning the concatenated multimodal inputs.
Lightweight Multimodal Fusion and Adaptation.
As LLMs become increasingly large, the computational costs for ERC also rise significantly Zhang et al. (2024b). Inspired by domain-specific LLM paradigms tailored for affective computing Hu et al. (2022); Li et al. (2023c); Tanioka et al. (2024), MSE-Adapter Yang et al. (2025) proposed a lightweight and adaptable plug-in architecture with two modules: TGM, for aligning textual and non-textual features, and MSF, for multi-scale cross-modal fusion. Built on a frozen LLM backbone and trained via backpropagation, it enables efficient and multimodally-aware ERC with minimal computational cost. Similarly, SpeechCueLLM Wu et al. (2025) introduced a lightweight plug-in that converts speech features into natural language prompts, enabling LLMs to perform multimodal emotion recognition without architectural changes.
@@ 6 Challenges and Prospects
FAIR-related Issues Pose Challenges in MERC.
@@
The FAIR principles provide guidelines for improving the Findability, Accessibility, Interoperability, and Reusability of digital assets Wilkinson et al. (2016). Collecting large and diverse multimodal emotion data is costly and time-consuming; some large dialogue datasets (e.g., M3ED, Zhao et al., 2022) are still monolingual and domain-bound. These limitations directly conflict with the FAIR principles. Some ERC datasets lack rich metadata or persistent identifiers, undermining findability and interoperability. Others are subject to access restrictions or copyright constraints, while many adopt inconsistent labeling schemes that hinder reusability. Consequently, researchers often have to train on small or biased samples, which undermines generalization and the reuse of models.
Low-Resource, Multilingual, and Multicultural Settings.
@@
As described in the previous paragraph, most state-of-the-art MERC systems are trained on English-language datasets, which limits their global applicability. Although building a large-scale, diverse MER corpus is essential, it remains an obvious challenge as it requires expert annotation of data. The limited annotated data forces researchers to rely on transfer learning Ananthram et al. (2020), zero-shot Qi et al. (2021), or few-shot methods Yang et al. (2023). However, data scarcity and the high cost of emotion annotation continue to be major obstacles for MER in low-resource domains Hussain et al. (2025). Emotions are expressed differently across languages and cultures, further compounding the challenges of MER. Variations in emotional expression and interpretation due to cultural differences can lead to inconsistencies in labeling. Most existing corpora are culture-specific, limiting their generalizability. Although researchers have acknowledged this challenge A. and V. (2024); Ghaayathri Devi et al. (2024), MER systems aiming for global applicability must account for both linguistic diversity and culturally driven display rules.
Complexities of Fusion Strategies across Modalities.
@@
Multimodal fusion techniques include early fusion, mid-level fusion, late fusion, hybrid fusion, and others Atrey et al. (2010); Gandhi et al. (2023). A key challenge is that conversational signals, such as voice, facial expressions, and transcripts, are inherently asynchronous and occur at different time scales, making it difficult to align them at the utterance level. Emotions also depend on the context of preceding and subsequent conversation turns, so the model must capture temporal dynamics Wang et al. (2024). Previous studies have used recurrent or self-attention layers to model sequential context Houssein et al. (2024); Dutta et al. (2024), but long-range dependencies remain challenging to learn. How to balance and integrate contextual sentiment cue features with multimodal fusion features in decision-making, and how to determine which fusion strategies are most effective across different modalities, remain open and important research questions A.V. et al. (2024); Ramaswamy and Palaniswamy (2024).
Cross-Modal Alignment, Noise Modality, Missing Modality, and Modality Conflicts.
@@
Misaligned or inconsistent features can inhibit the ability of a model to fully utilize multimodal signals, affecting its robustness and generalization Ma et al. (2024b); Li and Tang (2024). Noise modality, missing modality, or imbalanced modality distributions may bias simple fusion strategies Mai et al. (2024); Zhang et al. (2024a). Even when all modalities are available, they may convey conflicting emotional signals, further complicating fusion and decision-making. Perceiving the uncertainty of different modalities for feature enhancement and resolving conflicts between modality features are important areas that need further exploration in MER research. Therefore, exploring cross-modal transfer and fusion to improve generalization in ERC has attracted the attention of more and more researchers Fan et al. (2024); Li et al. (2024a); Feng and Fan (2025). Some ERC methods incorporate variants of cross-modal attention Guo et al. (2024); N. and Patil (2020), graph-based fusion Li et al. (2023a); Hu et al. (2021b), or mutual learning to align features during training and improve cross-domain performance Lian et al. (2021).
Effective Modality Selection.
@@
Multimodal learning refers to the integration of information from various heterogeneous sources and aims to effectively leverage data from diverse modalities He et al. (2024); Tsai et al. (2024). In multimodal representation learning, not all modalities contribute equally to the task. Some modalities may introduce noise and need to be removed, while others may not be essential for the task at hand but could be indispensable for other subtasks. Existing research proposes modality selection algorithms that identify the contribution of each modality Marinov et al. (2023); Mai et al. (2024). However, selecting the most appropriate subset of modalities for a specific task remains one of the key challenges in multimodal learning.
Efficient Fine-tuning Approach Using Multimodal LLMs.
@@
Multimodal LLMs have brought major advances in enabling machines to learn across modalities. Some models are increasingly used in MERC, offering zero-shot or few-shot generalization across different modalities Li et al. (2024c); Yang et al. (2024); Bo-Hao et al. (2025). The use of LLMs in MERC opens up new possibilities for capturing deeper semantic and conversational cues beyond surface-level emotion signals. However, efficiently fine-tuning these models for emotion understanding still presents challenges, especially in low-resource and culturally diverse settings. Efficient adaptation of MLLMs to capture the nuance of emotions across diverse datasets, languages, or cross-cultural settings remains an open research frontier.
MERC Application.

### v2

@@ 1 Introduction
However, human emotions are typically conveyed through multiple modalities, including auditory (e.g., speech), visual (e.g., facial expressions, gestures), and linguistic (e.g., the semantic content conveyed by transcribed text). As a result, recent research (e.g., Ma et al., 2024a; Van et al., 2025; Dutta and Ganapathy, 2025) has increasingly focused on multimodal settings in dialogue, a direction we refer to as Multimodal Emotion Recognition in Conversations (MERC). Researchers aim to identify the emotional state of a given utterance by integrating contextual information from different modalities, which often includes subtle personal emotional states such as happiness, anger, and hatred Poria et al. (2019b); Gong et al. (2024), thereby improving the effectiveness of emotion recognition in dialogues. Figure 1 illustrates an example of ERC with textual, acoustic, and visual inputs.
@@ 2 Task Settings and Review Methodology
Modalities.
In the context of MERC, we define a modality as a distinct source or channel of information that conveys emotional or communicative signals. These modalities correspond to observable representations derived from different sensory inputs and are typically processed in separate feature spaces. The most common modalities include:
  • Textual modality: The transcribed representation of spoken utterances, capturing the semantic and syntactic content of language.
  • Acoustic modality: Prosodic and paralinguistic features extracted from speech, such as tone, pitch, and energy.
  • Visual modality: Non-verbal cues such as facial expressions, head movements, eye gaze, and gestures.
@@
where $\,\cdot\,$$\cdot\,;\cdot\,$] denotes combination of modalities.
Methodology for Literature Compilation:
Strategy. We conduct a comprehensive literature search using sources such as the ACL Anthology, Google Scholar, and general search engines (e.g., Google Chrome). Within the ACL Anthology, we focus on top venues including EMNLP, ACL, NAACL, and related workshops. [^1] We used keywords such as “multimodal emotion recognition”, “emotion recognition in conversation”, “multimodal affective computing”, “dialogue emotion recognition”, “MERC benchmark”, “context-aware emotion detection”, and “multimodal conversational modeling”.
@@ 4.2 Context Modeling
Situation-level.
The emotional state of a speaker is influenced not only by the semantic content of the current utterance but also by the surrounding contextual semantics. Therefore, existing approaches Hu et al. (2021a); Majumder et al. (2019) commonly employ specialized networks to model the sequential dependencies among utterances, aiming to more accurately capture the speaker’s temporal emotional dynamics. Given the textual feature $u_{i}\in\mathbb{R}^{d_{u}}$ for each utterance, the sequential context representation $c^{s}_{i}\in\mathbb{R}^{2d_{u}}$ is computed as follows:
@@
Speaker-level.
Speaker identity information typically exhibits temporal and relational properties of emotions, which can enhance the model’s ability to perceive speaker role information. Therefore, to more effectively learn and distinguish speaker-level contextual representations, many studies have further introduced speaker-related structured information on top of dialogue context modeling. Common approaches include using Speaker Embeddings Ma et al. (2024a); Shen et al. (2020) to explicitly differentiate between different speakers, or leveraging Graph Neural Networks Ai et al. (2025a); Van et al. (2025) to construct interaction graphs between speakers, thereby more comprehensively modeling the dependencies between them:
Speaker Embeddings: The speaker embedding $S_{i}$ is combined with the modality features (e.g., text, audio, or vision) to produce modality representations that are speaker- and context-aware.
@@
Graph Neural Networks: Speaker interactions can be further modeled by constructing a dialogue graph to capture inter-utterance and inter-speaker dependencies beyond sequential semantics. A dialogue graph is typically defined as $G=(V,E,W,R)$, where each node $v_{i}\in V$ corresponds to the $i$-th utterance, and its associated feature vector $\mathbf{c}_{i}^{s}$ is obtained from sequential modeling of contextual dependencies. Edges $e_{ij}\in E$ represent interaction links between utterances, with associated weights $\omega_{ij}\in W$ reflecting the interaction strength and types $r_{ij}\in R$ encoding speaker-related or structural relationships.
@@ 5 Methodology
It is worth noticing that some methods in MERC inherently involve multiple components (e.g., fusion modules within graph-based or generation-based frameworks). Our taxonomy is not intended to be strictly disjoint, but rather to organize the literature based on the core modeling paradigm or innovation focus of each method. The categorization of methods is thus as follows:
  • Graph-based: Methods are categorized as graph-based when the primary architecture centers around graph neural networks for modeling conversational structure, even if auxiliary modules (e.g., fusion layers) are integrated.
  • Fusion-based: Methods are grouped under fusion-based when their main contribution lies in the design of cross-modal interaction mechanisms, regardless of the backbone architecture (e.g., Transformer, LSTM).
  • Generation-based: Generation-based methods refer to recent approaches that leverage LLMs to generate predictions or intermediate reasoning, often using prompt engineering or instruction tuning, even if lightweight fusion components are present.
@@ 5.1 Graph-based Methods
Traditional Graph Neural Networks. Previous works such as bc-LSTM Poria et al. (2017) and ICON Hazarika et al. (2018) primarily relied on sequential approaches. DialogueGCN Ghosal et al. (2019) was the first to introduce GNNs into ERC, addressing the limitations of earlier sequence-based models like DialogueRNN Majumder et al. (2019) in capturing contextual dependencies. To effectively integrate information from different modalities, Hu et al. (2021b) constructed a graph structure that fuses multimodal features, enabling the model to capture inter-modal dependencies through graph convolutional networks and incorporating speaker information to enhance the representation of conversational semantics. Inspired by the application of graph convolutions in ERC, Li et al. (2024b) proposed the GSF module, which introduces an alternating graph convolution mechanism to hierarchically extract both cross-modal and intra-modal emotional information. Some studies further enhanced graph-based models with attention mechanisms for multimodal fusion; for example, Feng and Fan (2025) integrated a Cross-modal Attention Module to better fuse information from different modalities, while Nguyen et al. (2023) designed a cross-modal attention mechanism to explicitly model the heterogeneity between modalities.
@@
Hypergraph Neural Networks. Although traditional graph-based methods can capture long-range and multimodal contextual information, they are often challenged by missing modalities during conversations. Lian et al. (2023) tackled this issue by jointly optimizing classification and reconstruction tasks in an end-to-end manner to effectively model incomplete data. The related works Li et al. (2023b); Huang et al. (2024a) considered the limitations imposed by the pairwise relationships between GNNs nodes.Van et al. (2025) constructed a multimodal fusion graph and introduced Hypergraph Neural Networks Feng et al. (2019) to connect multiple modalities or utterance nodes simultaneously, thereby capturing more complex multivariate dependencies and high-order interactions in conversations, and enhancing the modeling of emotion propagation.
@@
Fourier Graph Neural Networks. Increasing the depth of GNN layers can lead to the over-smoothing problem Liu et al. (2022); Yi et al. (2023), which hampers the modeling of long-range semantic dependencies and complementary modality relations. To address this issue, GS-MMC Ai et al. (2025b) proposes a graph-based framework for multimodal consistency and complementarity learning. This method employs a Fourier graph operator to extract high- and low-frequency emotional signals from the frequency domain, capturing both local variations and global semantic trends. Additionally, a contrastive learning mechanism van den Oord et al. (2019) is designed to enhance the semantic consistency and complementarity of these signals in a self-supervised manner, thereby improving the model’s ability to recognize true emotional states.
Graph-based methods effectively capture long-range dependencies and speaker interactions by modeling utterances as nodes and relations as edges. In traditional NLP tasks, they are particularly effective for structured inputs such as token-level representations Zhang et al. (2023, 2024b). However, integrating heterogeneous multimodal signals into graph structures remains challenging, as naïve connections may introduce noise without proper modality alignment.
@@ 5.2 Fusion-based Methods
Equal Modality Weights. Equal interaction can fully utilize information from various modalities, preventing over-reliance on a single modality. Li et al. (2022) proposed achieving emotion recognition by equally integrating emotional vectors and sentence vectors from different modalities to form emotion capsules. Zhang and Li (2023) designed a local constraint module for modalities within the Transformer to promote modality interaction and incorporates a semantic graph to address the lack of semantic relationship information between utterances. Mao et al. (2021) constructed a hierarchical Transformer where each modality can flexibly switch between sequential and feedforward structures based on contextual information. Inspired by hierarchical modality interaction, Ma et al. (2024a) introduced a hierarchical gating Ma et al. (2019) fusion strategy into the Transformer architecture to enable fine-grained modality interaction and designs self-distillation Zhang et al. (2019) to further learn better modality representations.
Text-Dominant Modality. Some methods propose models based on primary–auxiliary modality collaboration, where auxiliary modalities are used to enhance the performance of the primary (textual) modality. Huang et al. (2024b) suggested that enhancing a text-dominant model with auxiliary modalities can improve performance. Zou et al. (2022) employed a Transformer architecture to design cross-modal attention for learning fusion relationships between different modalities, preserving the integrity of the primary modality’s features while enhancing the representation of weaker modality features. It also uses a two-stage emotional cue extractor to extract emotional evidence. Building on this, Zou et al. (2023) proposed using weaker modalities as multimodal prompts while performing deep emotional cue extraction for stronger modalities. The cue information is embedded into various attention layers of the Transformer to facilitate the fusion of information between the primary and auxiliary modalities. Zhu et al. (2024) introduced an asymmetric CMA-Transformer module for central and auxiliary modalities to obtain fused modality information and proposes a hierarchical distillation Yang et al. (2021) framework to perform coarse- and fine-grained distillation. This approach ensures the consistency of modality fusion information at different granularities.
Fusion-based methods focus on learning cross-modal interactions through attention mechanisms, with Transformer-based models achieving strong generalization. These approaches are efficient for tasks with well-aligned modality inputs but often overlook dialogue-level structures such as speaker dependencies. Compared to graph-based models, they emphasize modality-level fusion over relational reasoning.
@@ 5.3 Generation-based Methods
In recent years, pretrained LLMs have achieved remarkable success in natural language processing tasks Chu et al. (2024), demonstrating strong emergent capabilities Wei et al. (2022). However, despite their powerful general-purpose abilities, leveraging their full potential in specific sub-tasks still requires carefully crafted, high-quality prompts Wei et al. (2021) to bridge the gap in reasoning capabilities.As shown in Figure 5, researchers have proposed various model improvement strategies to effectively integrate contextual and multimodal information into LLMs while addressing their substantial computational resource demands.
@@
Instruction-Tuned with Speaker and Context Modeling. ERC tasks have predominantly relied on discriminative modeling frameworks. With the emergence of LLMs, InstructERC Lei et al. (2023) was the first to propose a generative framework for ERC. It introduces a simple yet effective retrieval-based prompting module that helps LLMs explicitly integrate multi-granularity supervisory signals from dialogues. Additionally, it incorporates an auxiliary emotion alignment task to better model the complex emotional transitions between interlocutors in conversations. Inspired by the integration of commonsense knowledge in COSMIC Ghosal et al. (2020), recent work Fu (2024); Fu et al. (2025a) designed a prompt generation approach based on dialogue history to elicit speaker-related commonsense using LLMs by injecting commonsense knowledge into ERC.
Behavior-Aware and Multimodal Instruction-Tuned. To address the lack of multimodal integration, Dutta and Ganapathy (2025) incorporated both acoustic and textual modalities. Considering that visual information may provide richer emotional cues, Zhang et al. (2025) constructed a high-quality instruction dataset using image and text data, and fine-tunes the model using Low-Rank Adaptation (LoRA) Song et al. (2024). Furthermore, Fu et al. (2025b) introduced a novel behavior-aware Multimodal LLM (MLLM)-based ERC framework. It consists of three core components: a video-derived behavior generation module, a behavior alignment and refinement module, and an instruction tuning module Wei et al. (2021). The first two modules enable the model to infer human behaviors from limited information, thereby enhancing its behavioral perception capability. The instruction tuning module improves the model’s emotion recognition performance by aligning and fine-tuning the concatenated multimodal inputs.
Lightweight Multimodal Fusion and Adaptation. As LLMs become increasingly large, the computational costs for ERC also rise significantly Zhang et al. (2025). Inspired by domain-specific LLM paradigms tailored for affective computing Hu et al. (2022); Li et al. (2023c); Tanioka et al. (2024), MSE-Adapter Yang et al. (2025) proposed a lightweight and adaptable plug-in architecture with two modules: TGM, for aligning textual and non-textual features, and MSF, for multi-scale cross-modal fusion. Built on a frozen LLM backbone and trained via backpropagation, it enables efficient and multimodally-aware ERC with minimal computational cost. Similarly, SpeechCueLLM Wu et al. (2025) introduced a lightweight plug-in that converts speech features into natural language prompts, enabling LLMs to perform multimodal emotion recognition without architectural changes.
Generation-based methods leverage LLMs to reformulate ERC as a text generation task, enabling flexible adaptation across datasets and domains. Their end-to-end nature simplifies input processing but limits fine-grained control over multimodal integration. In contrast to structured models, LLMs excel at scalability but require further refinement to model multimodal dependencies explicitly.
@@ 6 Challenges and Prospects
FAIR-related Issues Pose Challenges in MERC. The FAIR principles provide guidelines for improving the Findability, Accessibility, Interoperability, and Reusability of digital assets Wilkinson et al. (2016). Collecting large and diverse multimodal emotion data is costly and time-consuming; some large dialogue datasets (e.g., M3ED, Zhao et al., 2022) are still monolingual and domain-bound. These limitations directly conflict with the FAIR principles. Some ERC datasets lack rich metadata or persistent identifiers, undermining findability and interoperability. Others are subject to access restrictions or copyright constraints, while many adopt inconsistent labeling schemes that hinder reusability. Consequently, researchers often have to train on small or biased samples, which undermines generalization and the reuse of models. To address these issues, future work could prioritize the development of multilingual benchmark datasets with standardized metadata and open licensing, possibly through collaborative consortiums that align with FAIR principles.
@@
Low-Resource, Multilingual, and Multicultural Settings. As described in the previous paragraph, most state-of-the-art MERC systems are trained on English-language datasets, which limits their global applicability. Although building a large-scale, diverse MER corpus is essential, it remains an obvious challenge as it requires expert annotation of data. The limited annotated data forces researchers to rely on transfer learning Ananthram et al. (2020), zero-shot Qi et al. (2021), or few-shot methods Yang et al. (2023). However, data scarcity and the high cost of emotion annotation continue to be major obstacles for MER in low-resource domains Hussain et al. (2025). Emotions are expressed differently across languages and cultures, further compounding the challenges of MER. Variations in emotional expression and interpretation due to cultural differences can lead to inconsistencies in labeling. Most existing corpora are culture-specific, limiting their generalizability. Although researchers have acknowledged this challenge A. and V. (2024); Ghaayathri Devi et al. (2024), MER systems aiming for global applicability must account for both linguistic diversity and culturally driven display rules. Future work could explore culture-adaptive pretraining and cross-lingual transfer learning methods that embed culturally sensitive emotion semantics across languages.
@@
Complexities of Fusion Strategies across Modalities. Multimodal fusion techniques include early fusion, mid-level fusion, late fusion, hybrid fusion, and others Atrey et al. (2010); Gandhi et al. (2023). A key challenge is that conversational signals, such as voice, facial expressions, and transcripts, are inherently asynchronous and occur at different time scales, making it difficult to align them at the utterance level. Emotions also depend on the context of preceding and subsequent conversation turns, so the model must capture temporal dynamics Wang et al. (2024). Previous studies have used recurrent or self-attention layers to model sequential context Houssein et al. (2024); Dutta et al. (2024), but long-range dependencies remain challenging to learn. How to balance and integrate contextual sentiment cue features with multimodal fusion features in decision-making, and how to determine which fusion strategies are most effective across different modalities, remain open and important research questions A.V. et al. (2024); Ramaswamy and Palaniswamy (2024). Recent advances in adaptive fusion strategies and dynamic attention mechanisms show promise, and future methods could explore transformer-based fusion that dynamically reweights modalities based on conversational context.
@@
Cross-Modal Alignment, Noise Modality, Missing Modality, and Modality Conflicts. Misaligned or inconsistent features can inhibit the ability of a model to fully utilize multimodal signals, affecting its robustness and generalization Ma et al. (2024b); Li and Tang (2024). Noise modality, missing modality, or imbalanced modality distributions may bias simple fusion strategies Mai et al. (2024); Zhang et al. (2024a). Even when all modalities are available, they may convey conflicting emotional signals, further complicating fusion and decision-making. Perceiving the uncertainty of different modalities for feature enhancement and resolving conflicts between modality features are important areas that need further exploration in MER research. Therefore, exploring cross-modal transfer and fusion to improve generalization in ERC has attracted the attention of more and more researchers Fan et al. (2024); Li et al. (2024a); Feng and Fan (2025). Some ERC methods incorporate variants of cross-modal attention Guo et al. (2024); N. and Patil (2020), graph-based fusion Li et al. (2023a); Hu et al. (2021b), or mutual learning to align features during training and improve cross-domain performance Lian et al. (2021). Future research can further investigate robust training frameworks that include modality dropout, uncertainty-aware fusion, or reinforcement learning to selectively attend to trustworthy modalities.
@@
Effective Modality Selection. Multimodal learning refers to the integration of information from various heterogeneous sources and aims to effectively leverage data from diverse modalities He et al. (2024); Tsai et al. (2024). In multimodal representation learning, not all modalities contribute equally to the task. Some modalities may introduce noise and need to be removed, while others may not be essential for the task at hand but could be indispensable for other subtasks. Existing research proposes modality selection algorithms that identify the contribution of each modality Marinov et al. (2023); Mai et al. (2024). However, selecting the most appropriate subset of modalities for a task remains one of the key challenges in multimodal learning. An emerging direction is to integrate learnable modality gates or sparsity-inducing regularization into fusion models to automatically suppress uninformative modalities during training.
@@
Efficient Fine-tuning Approach Using Multimodal LLMs. Multimodal LLMs have brought major advances in enabling machines to learn across modalities. Some models are increasingly used in MERC, offering zero-shot or few-shot generalization across different modalities Li et al. (2024c); Yang et al. (2024); Bo-Hao et al. (2025). The use of LLMs in MERC opens up new possibilities for capturing deeper semantic and conversational cues beyond surface-level emotion signals. However, efficiently fine-tuning these models for emotion understanding still presents challenges, especially in low-resource and culturally diverse settings. Efficient adaptation of MLLMs to capture the nuance of emotions across diverse datasets, languages, or cross-cultural settings remains an open research frontier. Promising future directions include using adapter modules or low-rank fine-tuning techniques to adapt large models to specific emotion tasks with minimal data.